# SGFormer: Spherical Geometry Transformer for $360^{\circ}$ Depth Estimation

### v1

@@ 3.3. SPDecoder
Bifuse (Wang et al., 2020) 0.1704 0.1528 0.7272 0.2466 0.7680 0.9251 0.9731
SliceNet (Pintore et al., 2021) 0.1254 0.1035 0.5761 0.1898 0.8575 0.9640 0.9867
Yun et al. (Yun et al., 2022) 0.0907 0.0658 0.4701 0.1502 0.9131 0.9792 0.9924
Pano3D (Albanis et al., 2021) PanoFormer (Shen et al., 2022a) 0.0699 0.0494 0.4046 0.1282 0.9436 0.9847 0.9939
EGFormer (Yun et al., 2023) 0.0660 0.0428 0.3874 0.1194 0.9503 0.9877 0.9952
@@
3.4. Loss Function
For a fair comparison, we strictly follow EGFormer (Yun et al., 2023), using the scale-and-shift-invariant loss function (Ranftl et al., 2020). Specifically, we align the predicted depth $D_{p}$ and the ground truth depth $D_{g}$ with their scale (s) and shift (t) appropriately through a least-squares criterion:
(7) $s,t=\arg\min\limits_{s,t}\left(s\cdot D_{p}+t-D_{g}\right)$
@@ 3.4. Loss Function
(9) $m=\left(D_{g}^{\prime}>0\right)$
(10) $L_{pix}=\frac{1}{n}\sum_{i=1}^{n}\left[m\cdot\left|\left(D_{p}^{\prime}-D_{g}^{\prime}\right)\right|\right]$
@@
(11) $L_{grad}=\sum_{x,y}\left(\left|G_{x}\left(D_{p}^{\prime}-D_{g}^{\prime}\right)_{x,y}\right|+\left|G_{y}\left(D_{p}^{\prime}-D_{g}^{\prime}\right)_{x,y}\right|\right)$
@@
(12) $L_{total}=L_{pix}+0.5\cdot L_{grad}$
@@
Figure 9. Qualitative results on the Structured3D dataset (the top two lines), and Pano3D dataset (the bottom two lines). More results can be found in Supplementary Materials.
4. Experiments
@@
We strictly follow the training and evaluation protocol of EGformer (Yun et al., 2023) to evaluate our proposed SGFormer approach with two solid benchmarks. The implementation details are described as follows.
4.1. Dataset and Related Details
@@
The two datasets employed for evaluation: Structured3D (Zheng et al., 2020) and Pano3D (Albanis et al., 2021). Structured3D is a high-quality synthetic dataset designed specifically to support a variety of structured 3D modeling tasks, providing many realistic images with detailed 3D structure annotations. On the other hand, the Pano3D dataset serves as a new benchmark for spherical panorama depth estimation. It aims to comprehensively evaluate the performance of depth prediction models, including the accuracy, the clarity of edges, and the smoothness of the image.
@@
The reasons for not using other datasets are as follows: Firstly, the PanoSunCG dataset (Wang et al., 2018) is no longer publicly available. Secondly, the images in the 3D60 dataset (Zioulis et al., 2018) exhibit depth value leakage through their pixel brightness. Thirdly, Pano3D incorporates a rendered version of the Matterport3D (Chang et al., 2017), employing Pano3D and Matterport3D for assessment would result in unnecessary duplication. Lastly, in the Stanford2D3D dataset (Armeni et al., 2017), the top and bottom parts of equirectangular images are not rendered properly. However, our method requires converting the north and south poles to the equatorial region through BRP. More details are included in the Supplementary Materials.
@@
In our experimental settings, we perform all experiments using a single GTX 3090 GPU, where the batch size is set to 1. We choose the AdamW (Loshchilov and Hutter, 2017) as the optimizer. On the Structured3D dataset, the initial learning rate is set at $5\cdot 10^{-5}$, with the model being trained for 50 epochs. Subsequently, after merging the Pano3D and Structured3D datasets, the model continues to be further trained for 20 epochs., maintaining the same learning rate, and adopting a learning rate exponential decay strategy of 0.95 to optimize the training effect.
4.2. Comparison Results
@@
Metrics: To evaluate the model performance fairly, we select a series of standard evaluation metrics in EGFormer (Yun et al., 2023), including absolute relative error (Abs.rel), squared relative error (Sq.rel), root mean square linear error (RMS.lin), and root mean square log error (RMSlog). The lower the value of these indicators, the better the performance of the model. Additionally, we use relative accuracy ($\delta^{n}$) as a measure of performance, where n represents different threshold levels (i.e., $\delta<1.25^{n}$). The higher the value of this indicator, the higher the accuracy of the predictions.
@@
Quantitative Analysis: To validate the effectiveness of our approach, we compare our SGFormer model with the current state-of-the-art methods, including strategies for bi-projection fusion, and long-range dependencies. As shown in Table 1, our model ranks first in nearly all evaluation metrics, indicating superior overall performance. Specifically, on the one hand, compared with Bifuse (Wang et al., 2020) based on bi-projection fusion, our model has achieved a 53% improvement (Abs.rel) and 41% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, the performance is significantly improved by 66% (Abs.rel) and 51% (RMS.lin). On the other hand, compared with EGFormer (Yun et al., 2023) based on long-range dependencies, our model achieves an 11% improvement (Abs.rel) and 12% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, SGFormer achieves 12% (Abs.rel) and 9% (RMS.lin) gain. In quantitative comparison, we achieve significant performance improvements because we fully leverage the geometric priors of the sphere to relieve the panoramic distortion. The lower quality of depth estimation in SliceNet (Pintore et al., 2021) is mainly due to the detail loss when processing equirectangular images in the vertical direction.
@@
Qualitative Analysis: In Fig. 9, we present a qualitative comparison with EGFormer (Yun et al., 2023) and PanoFormer (Shen et al., 2022a). We can observe that our performance is superior in terms of regional structural details. Specifically, as demonstrated in the red boxes in Fig. 9, more accurate depths are recovered. We perform better on local details because we provide additional spherical geometric structures.
@@
Table 2. Ablation studies of individual components. Starting from a baseline approach without any improvements, we add components one at a time. We conduct a series of ablation studies on the S3D (Structured3D) and Pano3D datasets. Results in bold indicate the best results. (CLE: curve local embedding, GCPE: global conditional position embedding, CR: circular rotation, BRP: bipolar re-projection)
@@
Figure 10. Qualitative comparisons about individual components. The top row shows visual comparisons of depth estimation after adding different components, while the bottom row shows the visual comparisons of the corresponding error maps for the estimated depth maps. The middle two rows display close-up views of the highlighted areas in both the top and bottom rows respectively. As we add more components from left to right, the object boundaries become clearer, and the overall depth estimation becomes more accurate with lower errors. The trend of error changes can be directly observed from the error maps. a, b, c, d, e are the same as Table 2.
4.3. Ablation study
@@
We conduct a thorough validation of the key components of the SGFormer model under the same experimental conditions. As shown in Table 2, initially, we establish a baseline with a model that does not include any specific components, using a Resnet-34 as the encoder and a ViT-based (Dosovitskiy et al., 2020) decoder. Then we sequentially introduce core components on the Structured3D and Pano3D datasets to explore the effectiveness of our model.
@@
Effectiveness of CLE: From Table 2, we can observe that on the Structured3D dataset, there is a 6.70% improvement (Abs.rel) and 6.55% (RMS.lin). On the Pano3D dataset, the performance is further boosted by 1.67% (Abs Rel) and 5.78% (RMS.lin). This illustrates that by incorporating CLE, we introduce additional spherical surface distance, which aids the network in perceiving the distribution of distortion.
@@
Effectiveness of GCPE: The introduction of GCPE compensates for the model’s geometric details and spatial structure across different resolutions. Results in Table 2 show that the performance is significantly improved by 9.12% (Abs.rel) and 3.43% (RMS.lin) on the structured3D dataset. For the Pano3D dataset, there is an improvement of 8.33% (Abs.rel) and 1.05% (RMS.lin).
@@
Effectiveness of CR: The CR component is added to achieve cross-window feature interaction. Comparisons in Table 2 reveal an 8.51% improvement (Abs.rel) and 2.42% (RMS.lin) on the Structured3D dataset. The performance is further enhanced by 3.54% (Abs.rel) and 4.11% (RMS.lin) on the Pano3D dataset.
@@
Effectiveness of BRP: The BRP component is incorporated to reduce distortion by projecting the north and south polar regions near the equator. Results from Table 2 show that performance is improved by 2.80% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, the performance is further boosted by 6.87% (Abs.rel) and 2.27% (RMS.lin).
@@
Fig. 10 shows the results of a qualitative comparison. As we gradually add CLE, GCPE, CR, and BRP components from left to right, the output depth maps not only perform better in overall depth estimation but also capture local structural details more precisely. Furthermore, through the visualization comparison of error maps, we can clearly see that the estimated error is continuously decreasing (The deeper the red color, the higher the error. The deeper the blue color, the lower the error).
5. Conclusion
@@
In this paper, to remove the adverse impacts of distortion, we integrate three spherical geometric priors: equidistortion, continuity, and surface distance, to propose a spherical geometry transformer, named SGFormer. Specifically, we employ BRP to project the polar regions with severe distortions towards the equator regions to relieve the distortion. This strategy is parameter-free and efficient without additional branches and complex transformation costs. CR leverages the continuity of the sphere, facilitating cross-window feature interaction and enhancing the understanding of the global structure. CLE generates relative position embeddings based on the pixel points projected onto the spherical space, providing additional spherical surface distance to perceive distortion explicitly. Additionally, we introduce a query-based global conditional position embedding strategy, which can dynamically provide varying spatial structure and geometric local details across different resolutions. Experiments on two popular benchmarks demonstrate that our proposed SGFormer significantly outperforms the SoTA methods.

### v2

@@ III-C SPDecoder
Bifuse [1] 0.1704 0.1528 0.7272 0.2466 0.7680 0.9251 0.9731
SliceNet [12] 0.1254 0.1035 0.5761 0.1898 0.8575 0.9640 0.9867
Yun et al. [36] 0.0907 0.0658 0.4701 0.1502 0.9131 0.9792 0.9924
Pano3D [16] PanoFormer [13] 0.0699 0.0494 0.4046 0.1282 0.9436 0.9847 0.9939
EGFormer [2] 0.0660 0.0428 0.3874 0.1194 0.9503 0.9877 0.9952
@@
III-D Loss Function
Similar to the existing methods [2], the scale-and-shift-invariant loss [37] is introduced to eliminate the influence of depth scale and offset differences between different datasets. Specifically, we align the predicted depth $D_{p}$ and the ground truth depth $D_{g}$ with their scale (s) and shift (t) appropriately through a least-squares criterion:
$s,t=\arg\min\limits_{s,t}\left(s\cdot D_{p}+t-D_{g}\right)$ (7)
@@ III-D Loss Function
$m=\left(D_{g}^{\prime}>0\right)$ (9)
$L_{pix}=\frac{1}{n}\sum_{i=1}^{n}\left[m\cdot\left|\left(D_{p}^{\prime}-D_{g}^{\prime}\right)\right|\right]$ (10)
@@
$L_{grad}=\sum_{x,y}\left(\left|G_{x}\left(D_{p}^{\prime}-D_{g}^{\prime}\right)_{x,y}\right|+\left|G_{y}\left(D_{p}^{\prime}-D_{g}^{\prime}\right)_{x,y}\right|\right)$ (11)
@@
$L_{total}=L_{pix}+\omega\cdot L_{grad}$ (12)
we empirically choose $\omega$ as 0.5.
@@
Figure 8: Qualitative results on the Structured3D dataset (the top two lines), and Pano3D dataset (the bottom two lines).
IV Experiments
@@
We follow the training and evaluation protocol of EGformer [2] to evaluate our proposed SGFormer approach with two solid benchmarks. The implementation details are described as follows.
IV-A Dataset and Implementation Details
@@
Two datasets are employed in our evaluation: Structured3D [15] and Pano3D [16]. Structured3D is a high-quality synthetic dataset designed specifically to support a variety of structured 3D modeling tasks, providing many realistic images with detailed 3D structure annotations. On the other hand, the Pano3D dataset serves as a new benchmark for spherical panorama depth estimation. It aims to comprehensively evaluate the performance of depth prediction models, including the accuracy, the clarity of edges, and the smoothness of the image.
@@
The reasons for not using other datasets are as follows: Firstly, the PanoSunCG dataset [38] is no longer publicly available. Secondly, the images in the 3D60 dataset [39] exhibit depth value leakage through their pixel brightness. Thirdly, Pano3D incorporates a rendered version of the Matterport3D [40], employing Pano3D and Matterport3D for assessment would result in unnecessary duplication. Lastly, in the Stanford2D3D dataset [41], the top and bottom parts of equirectangular images are not rendered properly. However, our method requires converting the north and south poles to the equatorial region through bipolar re-projection (BRP).
@@
In our experimental settings, we perform all experiments using a single GTX 3090 GPU, where the batch size is set to 1. We choose the AdamW [42] as the optimizer. On the Structured3D dataset, the initial learning rate is set at $5e^{-5}$, with the model being trained for 50 epochs. Subsequently, after merging the Pano3D and Structured3D datasets, the model continues to be further trained for 20 epochs, maintaining the same learning rate and adopting a learning rate exponential decay strategy of 0.95 to optimize the training effect.
TABLE II: Quantitative comparison on the S2D3D (Stanford2D3D) and M3D (Matterport3D) datasets with the SoTA methods. Blue indicates that our method achieves the best performance. means that we modify the structure of the HRDFuse network to ensure a fair comparison.
Dataset Method Pub’Year Abs.rel $\downarrow$ Sq.rel $\downarrow$ RMSE $\downarrow$ $\delta^{1}$uparrow $\delta^{2}$uparrow $\delta^{3}$uparrow
EGFormer [2] ICCV’23 0.1528 0.1408 0.4974 0.8185 0.9338 0.9736
PanoFormer [13] ECCV’22 0.1122 0.0786 0.3945 0.8874 0.9584 0.9859
OmniFusion [11] CVPR’22 0.1154 0.0775 0.3809 0.8674 0.9603 0.9871
S2D3D [41] UniFuse [10] RAL’21 0.1124 0.0709 0.3555 0.8706 0.9704 0.9899
Elite360D [43] CVPR’24 0.1182 0.0728 0.3756 0.8872 0.9684 0.9892
SGFormer (ours) - 0.1040 0.0581 0.3406 0.8998 0.9693 0.9908
EGFormer [2] ICCV’23 0.1473 0.1517 0.6025 0.8158 0.9390 0.9735
PanoFormer [13] ECCV’22 0.1051 0.0966 0.4929 0.8908 0.9623 0.9831
BiFuse [1] CVPR’20 0.1126 0.0992 0.5027 0.8800 0.9613 0.9847
BiFuse++ [44] TPAMI’22 0.1123 0.0915 0.4853 0.8812 0.9656 0.9869
M3D [40] UniFuse [10] RAL’21 0.1144 0.0936 0.4835 0.8785 0.9659 0.9873
OmniFusion [11] CVPR’22 0.1161 0.1007 0.4931 0.8772 0.9615 0.9844
HRDFUSE [9] CVPR’23 0.1172 0.0971 0.5025 0.8674 0.9617 0.9849
Elite360D [43] CVPR’24 0.1115 0.0914 0.4875 0.8815 0.9646 0.9874
SGFormer (ours) - 0.1039 0.0865 0.4790 0.8946 0.9642 0.9859
IV-B Comparison Results
@@
Metrics: To evaluate the model performance fairly, we select a series of standard evaluation metrics in EGFormer [2], including absolute relative error (Abs.rel), squared relative error (Sq.rel), root mean square linear error (RMS.lin), and root mean square log error (RMSlog). The lower the value of these indicators, the better the pmodel’s performance. Additionally, we use relative accuracy ($\delta^{n}$) as a measure of performance, where n represents different threshold levels (i.e., $\delta<1.25^{n}$). The higher the value of this indicator, the higher the accuracy of the predictions.
@@
Quantitative Analysis: To validate the effectiveness of our approach, we compare our SGFormer model with the current state-of-the-art methods, including strategies for bi-projection fusion and long-range dependencies. As shown in Table I, our model ranks first in nearly all evaluation metrics, indicating superior overall performance. Specifically, on the one hand, compared with Bifuse [1] based on bi-projection fusion, our model has achieved a 53% improvement (Abs.rel) and 41% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, the performance is significantly improved by 66% (Abs.rel) and 51% (RMS.lin). On the other hand, compared with EGFormer [2] based on long-range dependencies, our model achieves an 11% improvement (Abs.rel) and 12% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, SGFormer achieves 12% (Abs.rel) and 9% (RMS.lin) gain. In quantitative comparison, we achieve significant performance improvements because we fully leverage the geometric priors of the sphere to relieve the panoramic distortion. The lower quality of depth estimation in SliceNet [12] is mainly due to the detail loss when processing equirectangular images in the vertical direction.
@@
Qualitative Analysis: In Fig. 8, we present a qualitative comparison with EGFormer [2] and PanoFormer [13]. We can observe that our performance is superior in terms of regional structural details. Specifically, as demonstrated in the red boxes in Fig. 8, more accurate depths are recovered. We perform better on local details because we provide additional spherical geometric structures.
More Comparison Results: We also compare our method with the latest methods from 2024 on additional datasets. For a fair comparison, we strictly follow the Elite360D [43]. The quantitative comparison is shown in Table II, we achieve better performance on the Stanford2D3D dataset [41] and Matterport3D dataset [40].
@@
TABLE III: Ablation studies of individual components. We begin with a basline model, adding each component sequentially. We conduct a series of ablation studies on the S3D (Structured3D) and Pano3D datasets. Results in bold indicate the best results. (CLE: curve local embedding, GCPE: global conditional position embedding, CR: circular rotation, BRP: bipolar re-projection)
@@
Figure 9: Qualitative comparisons about individual components. The top row shows visual comparisons of depth estimation after adding different components, while the bottom row presents the visual comparisons of the error maps corresponding to the estimated depth maps. The middle two rows display close-up views of the highlighted areas in the top and bottom rows, respectively. As we add more components from left to right, the object boundaries become clearer, and the overall depth estimation becomes more accurate with lower errors. The trend of error changes can be directly observed from the error maps. a, b, c, d, e are the same as Table III.
IV-C Ablation study
@@
We conduct a thorough validation of the key components of the SGFormer model under the same experimental conditions. As shown in Table III, initially, we establish a baseline with a model that does not include any specific components, using a Resnet-34 as the encoder and a ViT-based [29] decoder. Then we sequentially introduce core components on the Structured3D and Pano3D datasets to explore the effectiveness of our model.
IV-C1 Effectiveness of curve local embedding (CLE)
@@
From Table III, we can observe that on the Structured3D dataset, there is a 6.70% improvement (Abs.rel) and 6.55% (RMS.lin). On the Pano3D dataset, the performance is further boosted by 1.67% (Abs Rel) and 5.78% (RMS.lin). These results illustrate that by incorporating CLE, we introduce additional spherical surface distance, which aids the network in perceiving the distribution of distortion.
IV-C2 Effectiveness of global conditional position embedding (GCPE)
@@
The introduction of GCPE compensates for the model’s geometric details and spatial structure across different resolutions. Results in Table III show that the performance is significantly improved by 9.12% (Abs.rel) and 3.43% (RMS.lin) on the structured3D dataset. For the Pano3D dataset, there is an improvement of 8.33% (Abs.rel) and 1.05% (RMS.lin).
IV-C3 Effectiveness of circular rotation (CR)
@@
The CR component is added to achieve cross-window feature interaction. Comparisons in Table III reveal an 8.51% improvement (Abs.rel) and 2.42% (RMS.lin) on the Structured3D dataset. The performance is further enhanced by 3.54% (Abs.rel) and 4.11% (RMS.lin) on the Pano3D dataset.
IV-C4 Effectiveness of bipolar re-projection (BRP)
@@
The BRP component is incorporated to reduce distortion by projecting the north and south polar regions near the equator. Results from Table III show that performance is improved by 2.80% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, the performance is further boosted by 6.87% (Abs.rel) and 2.27% (RMS.lin).
@@
In Fig. 9, as we gradually add CLE, GCPE, CR, and BRP components from left to right, the output depth maps perform better in overall depth estimation and capture local structural details more precisely. Furthermore, through the visualization comparison of error maps, we can see that the estimated error is continuously decreasing (Darker shades of the red indicate a higher error, while darker shades of the blue signify a lower error.).
IV-C5 Backbones
Following the previous works[11, 43], we use ResNet-34 to extract features from the ERP images. In addition, we also replace the backbone (ResNet-34) with a more advanced transformer-based architecture (i.e., Swin-B) to verify the extensibility of our proposed SPDecoder. As shown in Table IV, the combination of the advanced backbone and our designs can obtain a significant improvement across all metrics. This improvement indicates that although the traditional self-attention mechanism still can not address the distortion issue [13], our distortion-aware designs can fully use the advantages of these advanced backbones’ large receptive field, delivering superior performance. Considering the computational cost, we employ ResNet-34 as the backbone to extract ERP features.
TABLE IV: Results of using different backbones
Dataset Backbone Abs.rel $\downarrow$ Sq.rel $\downarrow$ RMS.lin $\downarrow$ RMSlog $\downarrow$ $\delta^{1}$uparrow
Pano3D [16] ResNet-34 0.0583 0.0364 0.3537 0.1101 0.9613
Swin-B 0.0519 0.0270 0.3066 0.0953 0.9725
V Conclusion
@@
In this paper, we propose a spherical geometry transformer, SGFormer, integrating three geometric priors: equidistortion, continuity, and surface distance, to mitigate distortion effects. We employ bipolar re-projection (BRP) to project polar regions towards the equator. This strategy is parameter-free and efficient without additional branches and complex transformation costs. Circular rotation (CR) leverages sphere continuity to enhance cross-window feature interaction and understanding of global structure. Curve local embedding (CLE) uses relative position embeddings from pixels projected onto spherical space, explicitly perceiving distortion with additional surface distance. Additionally, we introduce a query-based global conditional position embedding strategy, which can dynamically provide varying spatial structure and geometric local details across different resolutions. Experiments on two benchmarks demonstrate that SGFormer significantly outperforms SoTA methods.